### v1

# MDP Geometry, Normalization and Value Free Solvers

Abstract
Markov Decision Process (MDP) is a common mathematical model for sequential decision-making problems. In this paper, we present a new geometric interpretation of MDP, which is useful for analyzing the dynamics of main MDP algorithms. Based on this interpretation, we demonstrate that MDPs can be split into equivalence classes with indistinguishable algorithm dynamics. The related normalization procedure allows for the design of a new class of MDP-solving algorithms that find optimal policies without computing policy values.
1 Introduction
Markov Decision Process (MDP) is a common mathematical model for sequential decision making. It was initially introduced in the late 1950s with main algorithms such as Value Iteration (Bellman, 1957) and Policy Iteration (Howard, 1960). The topic has been developed over the course of the next decades, and the most important results were summarized in 1990 by Puterman (1990). More recently, the rise of practical reinforcement learning has drawn new attention to MDP analyses, leading to several important results. For the Policy Iteration algorithm, polynomial convergence for a fixed discount factor has been shown Ye (2011); Hansen et al. (2013); Scherrer (2013), and exponential convergence in the general case Fearnley (2010); Hollanders et al. (2012, 2016). For Value Iteration, it has been shown that the convergence is not strongly polynomial Feinberg and Huang (2014) and, under the connectivity assumption, has an extra convergence rate in addition to the discount factor $\gamma$ Mustafin et al. (2024). On conceptual level, the Value Function Polytope Dadashi et al. (2019) was an important finding for general MDP analysis.
In this paper, we introduce a new geometric interpretation of MDP and show that classical MDP problems are equivalent to certain geometrical problems. We also demonstrate the dynamics of classical algorithms in the new geometric space and introduce a new algorithm to solve an MDP without computing policy values.
2 Basic MDP setting
We assume a standard MDP setting Sutton and Barto (2018); Puterman (2014). An MDP is defined as $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathrm{st},\mathcal{P},\mathcal{R},\gamma\rangle$, where $\mathcal{S}=\{1,2,\ldots,n\}$ is a finite set of $n$ states; $\mathcal{A}$ is a finite set of all possible actions over all states combined; $\mathrm{st}$ is the map from $\mathcal{A}$ to $\mathcal{S}$ which sends an action $a$ defined over a state $s$ to the state $s$; $\mathcal{P}$ represents transitions specified by the transition probabilities of each action $P(s^{\prime}|a)$; $\mathcal{R}$ represents rewards given for actions in $\mathcal{A}$; and $\gamma\in(0,1)$ is a discount factor. Let us reiterate that in this setting, we assume that each action is attributed to a particular state via the map $\mathrm{st}$ (i.e. if in two different states there is an option to choose two actions with the same rewards and the same transition probability distribution, we consider them as two different actions).
Having an MDP, an agent chooses a policy to interact with it, a map $\pi:\mathcal{S}\rightarrow\mathcal{A}$ such that $\mathrm{st}\circ\pi=\mathrm{id}_{\mathcal{S}}$. We consider only deterministic stationary policies, which are policies that choose a single action to perform in each state throughout the entire trajectory. If policy $\pi$ chooses action $a$ in a state $s$, we say that $a\in\pi$. Each policy can be defined by the actions it takes in all states, $\pi=\{a_{1},\dots,a_{n}\}$, $s\in\mathcal{S}$.
Given a policy, we can define its value over a state $s$ to be the expected infinite sum of discounted rewards over trajectories starting from the state $s$:
$V_{\pi}(s)=\operatorname{\mathbb{E}}\left[\sum_{t=1}^{\infty}\gamma^{t}r_{t}\right],$
where $r_{t}$ is the reward received at time step $t$. The vector $V_{\pi}$ will be the unique vector that satisfies the Bellman equation $T^{\pi}V_{\pi}=V_{\pi}$, where $T^{\pi}$ is the Bellman operator, defined as:
$(T^{\pi}V)(s)=r_{\pi}+\sum_{s^{\prime}}P(s^{\prime}|\pi(s))\gamma V(s^{\prime}).$
Evaluating the policy values of a given policy $\pi$ is an important problem in MDPs called the Policy Evaluation Problem. However, the main problem of interest given an MDP is to find the optimal policy $\pi^{*}$ such that:
$V_{\pi^{*}}(s)\geq V_{\pi}(s)\quad\forall\pi,s.$
The exact solution to this problem can be found using the Policy Iteration algorithm, but each iteration requires inverting a square matrix of size $n\times n$, which is computationally expensive. Instead, we might be interested in a non-exact solution, or an $\epsilon$-optimal policy $\pi^{\epsilon}$—a policy that is worse than the optimal one by no more than $\epsilon$ in every state:
$V_{\pi^{*}}(s)-V_{\pi^{\epsilon}}<\epsilon\quad\forall s.$
This task can be solved by the Value Iteration algorithm, which is guaranteed to output an $\epsilon$-optimal policy in a number of iterations dependent on $\epsilon$, with each iteration having a computational complexity of $\mathcal{O}(n|\mathcal{A}|)$.
3 Geometry of action space
Let’s take an action $a$ on the state $1$ participating in a policy $\pi$ with the values $V^{\pi}$. We denote the reward associated with this action as $r_{a}$ and the probabilities to move to a state $s$ as $p^{a}_{s}$. We know that the policy values satisfy the Bellman equation:
$V^{\pi}(1)=r^{a}+\gamma p^{a}_{1}V^{\pi}(1)+\gamma p^{a}_{2}V^{\pi}(2)+\dots+\gamma p^{a}_{n}V^{\pi}(n).$
We then perform a simple manipulation: we move the value $V^{\pi}(1)$ to the right-hand side of the equation, obtaining
$0=r^{a}+(\gamma p^{a}_{1}-1)V^{\pi}_{1}+\gamma p^{a}_{2}V^{\pi}(2)+\dots+\gamma p^{a}_{n}V^{\pi}(n).$ (1)
This new equation can be presented as the dot product of two vectors: $(r^{a},\gamma p^{a}_{1}-1,\gamma p^{a}_{2},\dots,\gamma p^{a}_{n})$ and $(1,V^{\pi}_{1},V^{\pi}_{2},\dots,V^{\pi}_{n})$. Note that the two vectors are logically independent: the former vector contains all the information related to the action $a$ — reward and transition probabilities — and can be constructed without defining a policy. Similarly, the latter vector consists of policy values and can be constructed without knowing the actions participating in the policy. We denote the first vector as $a_{+}$ and call it the action vector, assuming it is a row vector, and the second vector as $V^{\pi}_{+}$ and call it the policy vector, assuming it is a column vector. With this notation we can establish an important proposition:
Proposition 3.1.
For a policy $\pi$ and an action $a\in\pi$, the dot product of the corresponding action vector and policy vector is 00. In other words, the action vector $a_{+}$ is orthogonal to the policy vector $V^{\pi}_{+}$ if $a\in\pi$.
Consider the $(n+1)$-dimensional linear space where the action vectors lie, which we call the action space. Index the coordinates starting from 00, so that for an action vector $a_{+}$ would have its zeroth coordinate equal to the reward $r^{a}$. The possible locations of action vectors in this space are limited. First, the sum of all vector entries except the zeroth one is equal to $\gamma-1$, which means that all of them lie on a fixed $n$-dimensional affine hyperplane. Additionally, exactly one of the last $n$ entries of an action vector $a_{+}$ must be negative, specifically the one corresponding to the state $\mathrm{st}(a)$, and all other entries are non-negative. Therefore, the actions of an MDP lie in the designated pairwise non-intersecting polytope zones in the action space, with each state having its own zone. For a state $s$, any corresponding action vector $x$ must obey the following inequalities:
[…]
Figure 1 shows an example of an action space for a 2-state MDP. We suggest to visualize the reward, the zeroth coordinate in the action space, as the height function, as presented in the figure. We will refer to this choice of a height function later when we talk about horizontal and vertical directions or the relation of being higher or lower.
Figure 1: An example of the action space for 2-states MDP with 3 actions on each state. Vertical axe show action rewards, and two horizontal axis correspond to the second (axe $p_{1}$) and third (axe $p_{2}$) entries of action vector.
From the point of view of the action space, the policy vector $V^{\pi}_{+}$ defines the hyperplane of vectors orthogonal to $V^{\pi}_{+}$. Denote this hyperplane by $\mathcal{H}_{\pi}$. From Proposition 3.1 we get that for any action $a\in\pi$ in the policy $\pi$, the corresponding action vector $a_{+}$ lies in the hyperplane $\mathcal{H}_{\pi}$. Therefore, an alternative definition of $\mathcal{H}_{\pi}$ for a policy $\pi=\{a_{1},\ldots,a_{n}\}$ is the hyperplane passing through the action vectors ${a_{1}}_{+},\ldots,{a_{n}}_{+}$. This allows us to restate the Bellman equation in more geometric terms.
Proposition 3.2.
The problem of evaluating the values of a policy $\pi=\{a_{1},\dots,a_{n}\}$ (Policy Evaluation Problem) is equivalent to the problem of calculating the hyperplane coefficients of the $n$-dimensional linear span $\mathcal{H}_{\pi}=\langle{a_{1}}_{+},\ldots,{a_{n}}_{+}\rangle$ of the participating action vectors ${a_{1}}_{+},\ldots,{a_{n}}_{+}$.
Proof.
$\implies$: If the policy values are known, from Proposition 3.1 we know that the hyperplane with the coefficients $(1,V^{\pi}_{1},V^{\pi}_{2},\dots,V^{\pi}_{n})$ pass through the action vectors.
[…]
The important property of this interpretation is that the coefficients of the hyperplane $\mathcal{H}_{\pi}$ can be recovered in an intuitive geometric way, presented in Figure 2. Recall that for a state $s$ we have a designated zone for action vectors defined by Equation 2. Consider the following one-dimensional line in the boundary of the zone: $x_{1},\ldots,\widehat{x_{s}},\ldots,x_{n}=0;x_{s}=\gamma-1$. Denote this line by $L_{s}$. In fact, this line represents all vectors of the self-loop actions in state $s$.
Figure 2:
Theorem 3.3.
The values of a policy $\pi$ can be recovered as the rewards of the intersection points between $\mathcal{H}_{\pi}$ and the self-loop lines $L_{s}$, $s\in\mathcal{S}$ divided by $1-\gamma$.
Proof.
For each state $s$, take the point $l_{s}$ of the intersection between $L_{s}$ and $\mathcal{H}_{\pi}$. We know that the vector $l_{s}$ is orthogonal to the vector $V^{\pi}_{+}$:
$0=l_{s}V^{\pi}_{+}=(l_{s})_{0}+\sum_{s^{\prime}=1}^{n}(l_{s})_{s^{\prime}}V^{\pi}_{s^{\prime}}=(l_{s})_{0}+(\gamma-1)V^{\pi}_{s}\implies V^{\pi}_{s}=(l_{s})_{0}/(1-\gamma).$
Thus, the zeroth coordinate of $l_{s}$ corresponds to the policy value in state $s$.∎
Theorem 3.3 helps to view policies in the geometric terms in the action space: every policy can be thought of as a hyperplane in this space, and any non-vertical hyperplane (i.e., non-parallel to the reward axis) can be thought of as a policy substitute, but not necessarily produced by a set of available actions. Such hyperplanes, or sets of values which are not produced by the actions in the MDP, often appear in practice, for example, during the run of the Value Iteration algorithm. To underline their similarity to regular policies, we call them pseudo-policies.
4 Optimal policy in action space
Consider an action $b$ and a policy $\pi$ such that $b\notin\pi$, assuming without loss of generality that $b$ is an action on the state 1. What is the dot product of the action vector $b_{+}$ and the policy vector $V^{\pi}_{+}$? Mathematically, it is:
$b_{+}V^{\pi}_{+}=r^{b}+(\gamma p^{b}_{1}-1)V^{\pi}_{1}+\dots+\gamma p^{b}_{n}V^{\pi}_{n}=r^{b}+\gamma\sum_{i=1}^{n}p^{b}_{i}V^{\pi}_{i}-V^{\pi}_{1}.$
This quantity is often used in the reinforcement learning literature and is known as an advantage or an appeal of action $b$ with respect to policy $\pi$, denoted as $\textrm{adv}(b,\pi)$ (in this paper we stick to the term advantage). It has an intuitive meaning: the quantity shows the potential gain (or loss) in the infinite sum of discounted rewards if, instead of following the policy $\pi$ from state 1, the agent does the action $b$ and then follows $\pi$.
Advantages are key quantities in MDPs since they define the dynamics of the main MDP algorithms:
  • Given a policy $\pi_{t}$, Policy Iteration algorithm chooses the action $a_{s}$ in a state $s$ with the highest advantage among other actions in $s$.
  • Similarly, while choosing actions to update values, Value Iteration algorithm chooses actions with the highest advantage with respect to current values.
Now let’s establish the relation between an action advantage and a policy hyperplane in the action space. Since the advantage of an action $b$ with respect to a policy $\pi$ is the inner product of the vector $b_{+}$ with the vector $V^{\pi}_{+}$, which is orthogonal to the hyperplane $\mathcal{H}_{\pi}$, a positive advantage means that the vector $b_{+}$ lies above $\mathcal{H}_{\pi}$, and a negative advantage means that $b_{+}$ lies below $\mathcal{H}_{\pi}$. With this observation, we are ready to present the second important result of this paper.
Theorem 4.1.
The problem of MDP solving, or finding a stationary optimal policy $\pi^{*}$ in an MDP, is equivalent to the problem of constructing a hyperplane $\mathcal{H}$ such that at least one action from every state is incidental to it and no action lies above it.
Proof.
$\implies$: Suppose we know the optimal policy $\pi^{*}$. Every action has 00 or negative advantage with respect to it; therefore, none of the action vectors lie above the hyperplane $\mathcal{H}_{\pi^{*}}$. On the other hand, each state has the optimal action, and the corresponding action vector is incidental to $\mathcal{H}_{\pi^{*}}$.
[…]
$a_{i}=\operatorname*{arg\,max}_{b}\textrm{adv}(b,\pi).$
Thus, policy $\pi$ is stationary with respect to greedy upgrade and, consequently, optimal. ∎
The theorem is sufficient to design an algorithm that solves an MDP without ever computing policy values by fitting the highest hyperplane, but we propose a different value-free algorithm later. We state a few more facts without a proof:
  • So far, we have discussed only deterministic policies; however, now we can see that stochastic policies are no different. If we have a stochastic policy that chooses action $a_{1}$ in state $s$ with probability $p$ and action $a_{2}$ with probability $1-p$ in the same state, we can construct an auxiliary action $a^{\prime}$, which has an action vector
    $a_{+}^{\prime}=p{a_{1}}_{+}+(1-p){a_{2}}_{+}$
    and analyze policy $\pi$ as a deterministic policy with the action choice $a^{\prime}$ in state $s$. For every set of actions $(a^{1}_{s},\dots,a^{k}_{s})$ in state $s$, stochastic policies might choose any point in the convex hull of the corresponding action vectors. It implies that unique stochastic optimal policy is impossible, since the hyperplane corresponding to the optimal policy should be tangent to the convex hull of action vectors and stochastic policies crosses its interior.
  • As we have shown, advantages are key quantities in MDPs. In our analysis, we think of actions as points in a vector space and policies as hyperplanes. However, the dual interpretation (which we call policy-first geometric space), where policies are vectors and actions are planes, is also possible and helpful for MDP analysis.
5 MDP transformation and affine equivalence
If we try to analyze the dynamics of the policy iteration algorithm in a 2-state MDP (which we do later in Section 6.1 and Figure 3), we see that it is not very convenient to analyze when the rewards for two states are not comparable (e.g., big rewards in one state and small rewards in another state), making all policy hyperplanes steep. Can we modify the MDP to make it flatter? It turns out we can! In this section, we describe the transformation $\mathcal{L}$ of an MDP which changes the values of every policy in one particular state, while keeping all advantages unchanged.
Choose a state $s$ and a real number $\delta\in\mathbb{R}$ — the desired increment of the state value $V^{\pi}(s)$ of every policy $\pi$. We denote this transformation as $\mathcal{L}_{s}^{\delta}$. To perform it, we need to modify every reward in the MDP using the following rule:
  • For every action $a,\mathrm{st}(a)=s$ the reward $r^{a}$ should be modified as:
    $r^{a}:=r^{a}-\delta(\gamma p^{a}_{s}-1),$
  • For every action $b,\mathrm{st}(b)\neq s$ the reward $r^{b}$ should be modified as:
    $r^{b}:=r^{b}-\delta\gamma p^{b}_{s},$
where $p^{a}_{s}$ and $p^{b}_{s}$ are the probabilities of getting to state $s$ from the state of the action.
The geometric interpretation of $\mathcal{L}_{s}^{\delta}$ is as follows. We transform the whole action space with the linear transformation
$(x_{0},x_{1},\ldots,x_{n})\mapsto(x_{0}-\delta x_{s},x_{1},\ldots,x_{n}).$
This transformation maps the designated action zones for each state to itself. Any point on a line $L_{s^{\prime}}$, $s^{\prime}\neq s$, is fixed by this transformation, and any point on the line $L_{s}$ is lifted by $\delta(1-\gamma)$. Coupled with Theorem 3.3, we see that the values of any hyperplane $\mathcal{H}$ stay the same for $s^{\prime}\neq s$, and increases by $\delta$ on the state $s$.
Next, we prove that the advantage of any action with respect to any policy does not change after the transformation. The first case is $\mathrm{st}(a)=s$. Denote the original value vector of a policy as $V_{+}$, the modified value vector as $V^{\prime}_{+}$, the original action vector as $a_{+}$, and the modified action vector as $a^{\prime}_{+}$. Then
$a_{+}V_{+}=\,$ $r^{a}+(\gamma p^{a}_{s}-1)V^{s}_{+}+\gamma\sum_{s^{\prime}\neq s}p^{s^{\prime}}_{a}V^{s^{\prime}}_{+}=$
$r^{a}-\delta(\gamma p^{a}_{s}-1)+(\gamma p^{a}_{s}-1)(V^{s}_{+}+\delta)+\gamma\sum_{s^{\prime}\neq s}p^{s^{\prime}}_{a}V^{s^{\prime}}_{+}=a^{\prime}_{+}V^{\prime}_{+}.$
For action $\mathrm{b}\neq s$ the proof looks essentially the same. Denoting the corresponding action vectors as $b_{+}$ and $b_{+}^{\prime}$ we have:
$b_{+}V_{+}=\,$ $r^{b}+\gamma p^{b}_{s}V^{s}_{+}+\gamma\sum_{s^{\prime}\neq s}p^{s^{\prime}}_{b}V^{s^{\prime}}_{+}=$
$r^{b}-\delta\gamma p^{b}_{s}+\gamma p^{b}_{s}(V^{s}_{+}+\delta)+\gamma\sum_{s^{\prime}\neq s}p^{s^{\prime}}_{b}V^{s^{\prime}}_{+}=b_{+}^{\prime}V_{+}^{\prime}.$
The transformation $\mathcal{L}_{s}^{\delta}$ has a few important properties. For MDPs $\mathcal{M}$ and MDP $\mathcal{M}^{\prime}$, we write $\mathcal{M}^{\prime}=\mathcal{L}_{s}^{\delta}\mathcal{M}$ if $\mathcal{L}_{s}^{\delta}$ transforms $\mathcal{M}$ to $\mathcal{M}^{\prime}$. Note that $\mathcal{M}$ and $\mathcal{M}^{\prime}$ have the same sets of actions and, therefore, the policies are drawn from the same sets of actions. Then, we can make the following observations regarding the dynamics of the policy iteration and value iteration algorithms in MDPs $\mathcal{M}$ and $\mathcal{M}^{\prime}$:
  • Each advantage $\textrm{adv}(a,\pi)$ remains the same, which leads to the same action choices during the policy improvement steps of the policy iteration algorithm in $\mathcal{M}$ and $\mathcal{M}^{\prime}$. Therefore, the transformation $\mathcal{L}_{s}^{\delta}$ preserves the dynamics of the Policy Iteration algorithm.
  • Assume the starting values of the Value Iteration algorithm are $V_{0}$ in $\mathcal{M}$ and $V_{0}^{\prime}$ in $\mathcal{M}^{\prime}$, such that $V_{0}(s)+\delta=V_{0}^{\prime}(s)$. Then, due to preservation of the advantages, the choice of action remains the same at every step of the algorithm. At the same time, since we increase the value of state $s$ by $\delta$ for every pseudo-policy, the difference between the two value vectors remains unchanged. This, in turn, implies that $\mathcal{L}_{s}^{\delta}$ preserves the error vectors $e_{t}=V_{t}-V^{*}$, which describe the dynamics of the value iteration algorithm and define $\epsilon$-optimal policies. It also preserves the values of $V_{t+1}-V_{t}$, the vector used to compute the stopping criteria in the classical Value Iteration algorithm. All these properties together imply that the transformation $\mathcal{L}_{s}^{\delta}$ preserves the dynamics of the Value Iteration algorithm if the initial values are adjusted. We provide more comments on the Value Iteration dynamics in section 6.2.
The dynamics of the main algorithms imply a similarity between MDPs $\mathcal{M}$ and $\mathcal{M}^{\prime}$, based on which we give the following definition:
Definition 5.1.
MDPs $\mathcal{M}$ and $\mathcal{M}^{\prime}$ are affine equivalent if there exists a sequence of operations $\mathcal{L}_{s_{1}}^{\delta_{1}},\mathcal{L}_{s_{2}}^{\delta_{2}},\dots,\mathcal{L}_{s_{k}}^{\delta_{k}}$ such that
$\mathcal{M}^{\prime}=\mathcal{L}_{s_{1}}^{\delta_{1}}(\mathcal{L}_{s_{2}}^{\delta_{2}}(\dots(\mathcal{L}_{s_{k}}^{\delta_{k}}\mathcal{M})\dots)).$
Affine equivalence is a transitive relation on the set of MDPs. An affine equivalence class of MDPs is a maximal subset of MDPs in which every two MDPs $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are affine equivalent.
Choose some MDP $\mathcal{M}$ and compute its optimal values $(V^{*}(1),\dots,V^{*}(n))$. Apply the following sequence of transformations to $\mathcal{M}$:
$\mathcal{M}^{*}=(\mathcal{L}_{1}^{-V^{*}(1)}\circ\mathcal{L}_{2}^{-V^{*}(2)}\circ\dots\circ\mathcal{L}_{n}^{-V^{*}(n)})\mathcal{M}.$ (3)
The values of the optimal policy $\pi^{*}$ in $\mathcal{M}^{*}$ are 0 in all states. At the same time, since the transformations $\mathcal{L}$ preserve advantages, the advantages of the modified actions $a^{\prime}\in\pi^{*}$ remain 00, which implies that
adv($\textrm{adv}(a^{\prime},\pi^{*})=r^{a^{\prime}}+\sum_{i=1}^{n}\gamma p^{a}_{i}V^{\pi^{\prime}}=r^{a^{\prime}}=0,$
meaning that every action participating in the optimal policy in $\mathcal{M}^{*}$ has reward 00, while every action not participating in the optimal policy has a reward equal to its advantage with respect to the optimal policy, which is always negative. Geometrically, the optimal policy hyperplane $\mathcal{H}_{\pi^{*}}$ is given by the equation $x_{0}=0$, all the action vectors from the optimal policy lie on the hyperplane, while every action not participating in the optimal policy lie below it. In this normal form, the reward of an action becomes a single measure of the goodness of an action.
This picture is convenient to analyze, so we give this special kind of MDPs a definition:
Definition 5.2.
We call an MDP $\mathcal{M}^{*}$ normal if all values of the optimal policy of $\mathcal{M}^{*}$ are equal to zero.
The normalization $\mathcal{M}^{*}$ of an MDP $\mathcal{M}$ is the result of applying the sequence of transformations $\mathcal{L}_{1}^{-V^{*}(1)},\mathcal{L}_{2}^{-V^{*}(2)},\dots,\mathcal{L}_{n}^{-V^{*}(n)}$ to the MDP $\mathcal{M}$ with the optimal policy values $V^{*}$. The normalization is always a normal MDP.
There is exactly one normal MDP in every affine equivalence class. In the next section, we use the normal MDP to analyze the dynamics of the main RL algorithms.
6 Dynamics of main RL algorithms
In this section, we demonstrate the geometrical dynamics of the main MDP algorithms —Value Iteration and Policy Iteration —in the action space. We analyze the dynamics of the normalized MDP since every MDP can be transformed into its normal form while keeping the same affine equivalence class.
6.1 Policy Iteration
Figure 3: Illustration of Policy Iteration dynamics in 2-state MDP.
The example dynamics of the policy iteration algorithm is presented in Figure 3. Policy iteration consists of two steps applied sequentially: policy evaluation and policy improvement. As follows from Proposition 3.2, the policy evaluation step is identical to hyperplane construction. For the policy improvement step, at each iteration, it simply chooses the highest actions with respect to the current policy $\pi_{t}$. Therefore, with every iteration policy iteration algorithm gets a higher hyperplane until the optimal one is reached.
6.2 Value Iteration
Figure 4: Illustration of value iteration dynamics in 2 states MDP.
The dynamics of the Value Iteration algorithm is a little bit more complex and is presented in Figure 4. At each step, value iteration performs the following update:
$V_{t+1}(s)$ $=\max_{a}\left[r(s,a)+\gamma\sum_{s^{\prime}}P(s^{\prime}|s,a)V_{t}(s^{\prime})\right]$
$=V_{t}(s)+r(s,a^{*})+(\gamma P(s|s,a^{*})-1)V_{t}(s)+\gamma\sum_{s^{\prime}\neq s}P(s^{\prime}|s,a^{*})V_{t}(s^{\prime})$
$=V_{t}(s)+(1-\gamma)\textrm{adv}(a^{*},V_{t}).$
Having this interpretation, it is easy to graphically demonstrate an additional convergence factor described in Mustafin et al. (2024): assume that all current errors $e_{t}=V_{t}-V^{*}$ are positive (in a normalized MDP $e_{t}=V_{t}$). Let’s take a state $s$ with the highest error $e_{t}(s)$. The advantage of the action $a^{*}$ in state $s$, which maximizes advantage with respect to pseudo-policy $\pi_{t}$ (which has values $V_{t}$), is negative. At the same time, its absolute value would be exactly $(1-\gamma)V_{t}(s)$ only if $a^{*}$ is an action participating in the optimal policy and lying exactly on the edge, i.e. this action is a self-loop action. If $a^{*}$ does not participate in the optimal policy, its advantage is at least $(1-\gamma)a^{*}_{+}V{t+}+\delta^{\prime}$, where $\delta^{\prime}$ is the minimum disadvantage of a non-optimal action. If $a^{*}$ does participate in the optimal policy, but it is not a self-loop action, it lies “inside” its state zone and the absolute value of its advantage is again higher than $(1-\gamma)a^{*}+{V_{t}}_{+}$. These two factors guarantee geometric convergence with a rate faster than $\gamma$.
Geometrically, each iteration of the Value Iteration algorithm not only brings the pseudo-policy implied by the current values closer to optimal policy, but also skews its slope closer to a horizontal. This extra transformation enables achieving $\epsilon$-optimal policy when errors $e_{t}=V_{t}-V^{*}$ are still significantly high. Please refer to Figure 4 and the original paper (Mustafin et al., 2024) for details.
7 Value-free solvers
There is one more nice property of a normalized form of a MDP that we did not mention in Section 5: it is trivially solvable! To solve it, one just needs to pick the action with a reward of 0 at every state, and the resulting policy is optimal. This policy has a value of 0 at every state, while any other action has a negative reward, thus their inclusion to a policy will result in negative values.
This is an intriguing property of an MDP. Imagine one needs to solve an MDP $\mathcal{M}$. Suppose that $\mathcal{M}$ can be solved by Policy Iteration starting from a policy $\pi_{0}$ in $T$ steps. If the same algorithm is run on the normalized MDP $\mathcal{M}^{*}$, it will take exactly the same number of iterations $\pi_{0}$ to solve it as it is shown in Section 6.1. However, there is no need to run the algorithm on the normalized MDP: we can just pick the optimal actions with rewards 00 to form the optimal policy. This property of the normalized MDP gives an idea for a new solving algorithm.
Given a random MDP $\mathcal{M}$, it is not possible to normalize it without solving it, because the optimal values are not known. However, we still have the transformations $\mathcal{L}_{s}^{\delta}$ available to us, and we have a goal: if at some point we achieve an MDP in which the maximum action reward in every state is 0, we know that the corresponding actions form the optimal policy. Furthermore, the following lemma sets a stopping criteria for an approximate solution.
Lemma 7.1.
Take the policy $\pi$ in an MDP $\mathcal{M}$ that chooses the actions with the maximum rewards in every state. If the overall maximum reward is 0, and the minimum reward of the actions in the policy $\pi$ is $r_{min}<0$, the policy is $\epsilon$-optimal with $\epsilon=-r_{min}/(1-\gamma)$.
Proof.
We know that values of any policy in this MDP are upper bounded by 0. At the same time, minimum possible values of the policy satisfy $||V^{*}-V^{\pi}||_{\infty}<(0-r_{min})/(1-\gamma)$. ∎
Having established the stopping criteria, we are now ready to provide an intuition for the solving algorithm. For simplicity, we start by decreasing the reward of every action in the MDP by the overall maximum reward such that the maximum action reward becomes 00 and every action has a non-positive reward. Then our goal is to carefully “flatten” action rewards by applying transformations $\mathcal{L}_{s}^{\delta}$ to the states where the highest rewards are below 0, so that the maximum reward in every state becomes closer to 00. The problem is that while increasing the reward of an action $a$ in a state $s$, the transformation $\mathcal{L}_{s,\delta}$ also has a negative side effect: it decreases the rewards of all actions in every other state that have a non-zero transition probability to $s$. The straightforward approach would be to take these side effects into account, but in this case, computing the correct values $\delta_{1},\dots,\delta_{n}$ would require inverting the probability matrix (in fact, this algorithm is identical to Policy Iteration). The easiest way to simplify the algorithm is to not take these “side” effects into account, which leads to Algorithm 1.
Algorithm 1 Value Free approximate solver (VFS)
  Initialize $\mathcal{M}_{0}:=\mathcal{M}$ and $t=1$.
  Update:
  • $\forall s$ compute $\delta_{s}=\min_{a}r_{a}/(\gamma p^{a}_{s}-1)$.
  • Update the MDP $\mathcal{M}_{t}=(\mathcal{L}_{1}^{\delta_{1}}\circ\dots\circ\mathcal{L}_{n}^{\delta_{n}})\mathcal{M}_{t-1}$.
  • Store the maximum action reward in each state while performing the update. Set $R^{m}_{t}$ to the minimum of these maxima.
  if  $|R^{m}_{t}|/(1-\gamma)\geq\epsilon$ then
     Increment $t$ by 1 and return to the update step.
  else
     Output policy $\pi$, such that $\pi(s)=\operatorname*{arg\,max}_{a}r(a)$.
  end if
Algorithm 1 has the same iteration complexity as the Value Iteration algorithm but has different properties. First, Value Iteration convergence depends on the initial value guess, while VFS performance does not depend on the quality of initialization. Second, as we know from Mustafin et al. (2024), the Value Iteration algorithm benefits from information exchange between the nodes. In particular, it does not perform well on tree-structured MDPs, potentially achieving the worst possible convergence rate of $\gamma$. In contrast, tree-structured MDPs are the best case for the VFS algorithm. We can show this for a broader case of hierarchical MDPs: MDPs in which every state can be attributed to a class $c\in\mathbb{N}$ such that every action leads either back to the state $s$ or to states with a lower class. We can observe that after one step of the algorithm, every node of class 1—leaf nodes—will have a maximum reward of 0. Then, since there are no states downstream, these states remain unchanged during the run of the algorithm, which, in turn, means they won’t affect the nodes of higher classes. This observation allows a simple induction proof of the fact that VFS will converge in $C$ iterations in a hierarchical MDP, where $C$ is the number of hierarchy classes.
These two properties suggest that while having the same iteration computational complexity, VFS and VI are two different algorithms. However, it is possible that VFS is equivalent to VI with some initialization strategy and perhaps some modification of the MDP (for example, removing self-loops). The next section demonstrates the experimental comparison of the two algorithms.
8 Experimental results
In the section we present the performance results of the VFS Algorithm 1 on three experiments. We compare the convergence of the VFS algorithm with the Value Iteration algorithm with different learning rates. The comparison is done on three kinds of MDPs:
  • Random MDP: MDP with 10 states, 1 to 3 actions on every state, and a random reward.
  • Grid World MDP: 100 states correspond to the cells in a 10 by 10 grid. Possible actions are “Up”, “Left”, “Down” and “Right”, which are available at any state unless they bring the agent behind the border of the grid. Rewards for each action are equal to the sum of the cell coordinates plus a small random component.
  • Cycle MDP: numbered states organized into a cycle. Each state has 3 available actions, which move the agent to the first, second and third state forward. Reward is equal to the number of the state plus small random component.
Additionally to the parameters above, all the MDPs are created with an extra “execution probability” parameter. Each action’s transition probabilities are multiplied by the execution probability, and the remainder probability is assigned as the probability to stay in the same state (self-loop).
Our results are presented in Figure 5. Subfigures show that when execution probability is set to $1$, VI and VFS algorithms have similar performance, but then as we decrease it, the performance of VFS improves and the performance of the VI declines.
(a) Random MDP
(b) Grid World
(c) Ring
(d) Legend
Figure 5: Experimental results for 3 kind of MDPs. X-axis: action execution probability. Y-axis: average number of algorithm iterations before convergence to $\epsilon$. Green colors - VFS algorithm, Red colors- VI algorithms. Circle marker with light color - $0.5$ learning rate, Square marker with medium color - $0.75$ learning rate, Upper triangle marker with dark color - $1.0$ learning rate.
9 Discussion and Concluding remarks
In this paper, we have presented a new geometric interpretation of Markov Decision Processes, establishing the equivalence between MDP problems and particular geometrical problems. Based on this interpretation, we developed a new MDP solving algorithm that does not explicitly compute policy values to solve the MDP. At the same time, this paper does not provide final results but rather invites researchers in the field to apply the presented new interpretation to MDP problems. The authors see a few potential future research directions:
  • The new interpretation might be used to analyze the convergence of existing MDP algorithms.
  • The convergence of the presented value-free algorithm remains an open question, as its equivalence to the Value Iteration algorithm remains unresolved.
  • The value-free algorithm presented in the paper works only in a known MDP setting. However, it might be possible to develop a version that can be applied in a learning setting, possibly with deep neural networks (similarly to how DQN is based on classical Q-learning).