### v2

# MDP Geometry, Normalization and Value Free Solvers

Abstract
The Markov Decision Process (MDP) is a widely used mathematical model for sequential decision-making problems. In this paper, we present a new geometric interpretation of MDPs. Based on this interpretation, we show that MDPs can be divided into equivalence classes with indistinguishable key solving algorithms dynamics. This related normalization procedure enables the development of a novel class of MDP-solving algorithms that find optimal policies without computing policy values. The new algorithms we propose for different settings achieve and, in some cases, improve upon state-of-the-art results.
1 Introduction
Markov Decision Process (MDP) is a common mathematical model for sequential decision making. It was initially introduced in the late 1950s with main algorithms such as Value Iteration (Bellman, 1957) and Policy Iteration (Howard, 1960). The topic has been developed over the course of the next decades, and the most important results were summarized in 1990 by Puterman (1990). More recently, the rise of practical reinforcement learning has drawn new attention to MDP analyses, leading to several important results. For the Policy Iteration algorithm, polynomial convergence for a fixed discount factor has been shown Ye (2011); Hansen et al. (2013); Scherrer (2013), and exponential convergence in the general case Fearnley (2010); Hollanders et al. (2012, 2016). For Value Iteration, it has been shown that the convergence is not strongly polynomial Feinberg and Huang (2014) and, under the connectivity assumption, has an extra convergence rate in addition to the discount factor $\gamma$ Mustafin et al. (2024). On conceptual level, the Value Function Polytope Dadashi et al. (2019) was an important finding for general MDP analysis.
In this paper, we introduce a new geometric interpretation of an MDP and show how this interpretation helps to analyze old and develop new algorithms to solve MDPs. Out contribution is as follows:
  • We introduce a new geometric interpretation of the classical MDP setting and show that the common MDP problems of policy evaluation and finding an optimal policy have geometric analogies (Sections 3, 4).
  • We demonstrate that there is a transformation that modifies the MDP while preserving the dynamics of the main solver algorithms. Based on this transformation, we define an affine equivalence relation and an affine equivalence class. We show that in every class, there are normal and near-normal MDPs for which it is trivial to find approximate and exact solutions (Sections 5).
  • Based on this intuition, we propose a new value-free MDP-solving algorithm — Reward Balancing — which solves an MDP by modifying it such that it remains in the same affine equivalence class while approaching normal form. We show that this algorithm, while having the same iteration complexity as Value Iteration, exhibits different convergence properties. We identify the cases where the new algorithm converges faster compared to classical VI (Section 7).
  • We extend the Reward Balancing algorithm to the unknown MDP case (similar to Q-learning). We demonstrate that this algorithm achieves the state-of-the-art theoretical convergence rate while being fully parallelizable, thereby improving upon state-of-the-art results for parallel computation in the stochastic MDP case (Section 8).
2 Basic MDP setting
We assume a standard MDP setting Sutton and Barto (2018); Puterman (2014). An MDP is defined as $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathrm{st},\mathcal{P},\mathcal{R},\gamma\rangle$, where $\mathcal{S}=\{s_{1},s_{2},\ldots,s_{n}\}$ is a finite set of $n$ states; $\mathcal{A}$ is a finite set of size $m$ of all possible actions over all states combined; $\mathrm{st}$ is the map from $\mathcal{A}$ to $\mathcal{S}$ which sends an action $a$ defined over a state $s$ to the state $s$; $\mathcal{P}$ represents transitions specified by the transition probabilities of each action $P(s^{\prime}|a)$; $\mathcal{R}$ represents rewards given for actions in $\mathcal{A}$; and $\gamma\in(0,1)$ is a discount factor. Let us reiterate that in this setting, we assume that each action is attributed to a particular state via the map $\mathrm{st}$ (i.e. if in two different states there is an option to choose two actions with the same rewards and the same transition probability distribution, we consider them as two different actions).
Having an MDP, an agent chooses a policy to interact with it, a map $\pi:\mathcal{S}\rightarrow\mathcal{A}$ such that $\mathrm{st}\circ\pi=\mathrm{id}_{\mathcal{S}}$. We consider only deterministic stationary policies, which are policies that choose a single action to perform in each state throughout the entire trajectory. If policy $\pi$ chooses action $a$ in a state $s$, we say that $a\in\pi$. Each policy can be defined by the actions it takes in all states, $\pi=\{a_{1},\dots,a_{n}\}$, $s\in\mathcal{S}$, where $n$, again, is the number of states.
Given a policy, we can define its value over a state $s$ to be the expected infinite sum of discounted rewards over trajectories starting from the state $s$:
$V_{\pi}(s)=\operatorname{\mathbb{E}}\left[\sum_{t=1}^{\infty}\gamma^{t}r_{t}\right],$
where $r_{t}$ is the reward received at time step $t$ starting with state $s$ and actions taken from policy $\pi$. The vector $V_{\pi}$ will be the unique vector that satisfies the Bellman equation $T^{\pi}V_{\pi}=V_{\pi}$, where $T^{\pi}$ is the Bellman operator, defined as:
$(T^{\pi}V)(s)=r_{\pi}+\sum_{s^{\prime}}P(s^{\prime}|\pi(s))\gamma V(s^{\prime}).$
Evaluating the policy values of a given policy $\pi$ is an important problem in MDPs called the Policy Evaluation Problem. However, the main problem of interest given an MDP is to find the optimal policy $\pi^{*}$ such that:
$V_{\pi^{*}}(s)\geq V_{\pi}(s)\quad\forall\pi,s.$
The exact solution to this problem can be found using the Policy Iteration algorithm, but each iteration requires inverting a square matrix of size $n\times n$, which is computationally expensive. Instead, we might be interested in a non-exact solution, or an $\epsilon$-optimal policy $\pi^{\epsilon}$—a policy that is worse than the optimal one by no more than $\epsilon$ in every state:
$V_{\pi^{*}}(s)-V_{\pi^{\epsilon}}<\epsilon\quad\forall s.$
This task can be solved by the Value Iteration algorithm, which is guaranteed to output an $\epsilon$-optimal policy in a number of iterations dependent on $\epsilon$, with each iteration having a computational complexity of $\mathcal{O}(n|\mathcal{A}|)$.
3 Geometry of action space
Let’s take an action $a$ on the state $1$ participating in a policy $\pi$ with the values $V^{\pi}$. We denote the reward associated with this action as $r_{a}$ and the probabilities to move to a state $s$ as $p^{a}_{s}$. We know that the policy values satisfy the Bellman equation:
$V^{\pi}(1)=r^{a}+\gamma p^{a}_{1}V^{\pi}(1)+\gamma p^{a}_{2}V^{\pi}(2)+\dots+\gamma p^{a}_{n}V^{\pi}(n).$
We then perform a simple manipulation: we move the value $V^{\pi}(1)$ to the right-hand side of the equation, obtaining
$0=r^{a}+(\gamma p^{a}_{1}-1)V^{\pi}_{1}+\gamma p^{a}_{2}V^{\pi}(2)+\dots+\gamma p^{a}_{n}V^{\pi}(n).$ (1)
This new equation can be presented as the dot product of two vectors: $(r^{a},\gamma p^{a}_{1}-1,\gamma p^{a}_{2},\dots,\gamma p^{a}_{n})$ and $(1,V^{\pi}_{1},V^{\pi}_{2},\dots,V^{\pi}_{n})$. Note that the two vectors are logically independent: the former vector contains all the information related to the action $a$ — reward and transition probabilities — and can be constructed without defining a policy. Similarly, the latter vector consists of policy values and can be constructed without knowing the actions participating in the policy. We denote the first vector as $a_{+}$ and call it the action vector, assuming it is a row vector, and the second vector as $V^{\pi}_{+}$ and call it the policy vector, assuming it is a column vector. With this notation we can establish the following proposition:
Proposition 3.1.
For a policy $\pi$ and an action $a\in\pi$, the dot product of the corresponding action vector and policy vector is 00. In other words, the action vector $a_{+}$ is orthogonal to the policy vector $V^{\pi}_{+}$ if $a\in\pi$.
Consider the $(n+1)$-dimensional linear space where the action vectors lie, which we call the action space. Index the coordinates starting from 00, so that for an action vector $a_{+}$ would have its zeroth coordinate equal to the reward $r^{a}$. The possible locations of action vectors in this space are limited. First, the sum of all vector entries except the zeroth one is equal to $\gamma-1$, which means that all of them lie on a fixed $n$-dimensional affine hyperplane. Additionally, exactly one of the last $n$ entries of an action vector $a_{+}$ must be negative, specifically the one corresponding to the state $\mathrm{st}(a)$, and all other entries are non-negative. Therefore, the actions of an MDP lie in the designated pairwise non-intersecting polytope zones in the action space, with each state having its own zone. For a state $s$, any corresponding action vector $x$ must obey the following inequalities:
[…]
Figure 1 shows an example of an action space for a 2-state MDP. We suggest to visualize the reward, the zeroth coordinate in the action space, as the height function, as presented in the figure. We will refer to this choice of a height function later when we talk about horizontal and vertical directions or the relation of being higher or lower.
Figure 1: An example of the action space for 2-states MDP with 3 actions on each state. Vertical axe show action rewards, and two horizontal axis correspond to the second (axe $p_{1}$) and third (axe $p_{2}$) entries of action vector.
From the point of view of the action space, the policy vector $V^{\pi}_{+}$ defines the hyperplane of vectors orthogonal to $V^{\pi}_{+}$. Denote this hyperplane by $\mathcal{H}_{\pi}$. From Proposition 3.1 we get that for any action $a\in\pi$ in the policy $\pi$, the corresponding action vector $a_{+}$ lies in the hyperplane $\mathcal{H}_{\pi}$. Therefore, an alternative definition of $\mathcal{H}_{\pi}$ for a policy $\pi=\{a_{1},\ldots,a_{n}\}$ is the hyperplane passing through the action vectors ${a_{1}}_{+},\ldots,{a_{n}}_{+}$. This allows us to restate the Bellman equation in more geometric terms.
Proposition 3.2.
The problem of evaluating the values of a policy $\pi=\{a_{1},\dots,a_{n}\}$ (Policy Evaluation Problem) is equivalent to the problem of calculating the coordinates of a normal vector to the $n$-dimensional linear span $\mathcal{H}_{\pi}=\langle{a_{1}}_{+},\ldots,{a_{n}}_{+}\rangle$ of the participating action vectors ${a_{1}}_{+},\ldots,{a_{n}}_{+}$.
Proof.
$\implies$: If the policy values are known, from Proposition 3.1 we know that the hyperplane with the coefficients $(1,V^{\pi}_{1},V^{\pi}_{2},\dots,V^{\pi}_{n})$ pass through the action vectors.
[…]
The important property of this interpretation is that the values of the policy $\pi$ corresponded to the hyperplane $\mathcal{H}_{\pi}$ can be recovered in an intuitive geometric way, presented in Figure 2. Recall that for a state $s$ we have a designated zone for action vectors defined by Equation 2. Consider the following one-dimensional line in the boundary of the zone: $x_{1},\ldots,\widehat{x_{s}},\ldots,x_{n}=0;x_{s}=\gamma-1$. Denote this line by $L_{s}$. In fact, this line represents all vectors of the self-loop actions in state $s$.
Figure 2: An example of action states for a 2-state MDP with 3 actions on each state, with the policy hyperplane $\mathcal{\pi}$ of policy $\pi=(a,b)$ added. The black line connecting two actions represents the intersection of the policy hyperplane and the constraint hyperplane. The two red bars represent the values of the policy $\pi$ in states 1 and 2.
Theorem 3.3.
The values of a policy $\pi$ can be recovered as the rewards of the intersection points between $\mathcal{H}_{\pi}$ and the self-loop lines $L_{s}$, $s\in\mathcal{S}$ divided by $1-\gamma$.
Proof.
For each state $s$, take the point $l_{s}$ of the intersection between $L_{s}$ and $\mathcal{H}_{\pi}$. We know that the vector $l_{s}$ is orthogonal to the vector $V^{\pi}_{+}$:
$0=l_{s}V^{\pi}_{+}=(l_{s})_{0}+\sum_{s^{\prime}=1}^{n}(l_{s})_{s^{\prime}}V^{\pi}_{s^{\prime}}=(l_{s})_{0}+(\gamma-1)V^{\pi}_{s}\implies V^{\pi}_{s}=(l_{s})_{0}/(1-\gamma).$
Thus, the zeroth coordinate of $l_{s}$ corresponds to the policy value in state $s$.∎
Theorem 3.3 helps to view policies in the geometric terms in the action space: every policy can be thought of as a hyperplane in this space, and any non-vertical hyperplane (i.e., non-parallel to the reward axis) can be thought of as a policy substitute, but not necessarily produced by a set of available actions. Such hyperplanes, or sets of values which are not produced by the actions in the MDP, often appear in practice, for example, during the run of the Value Iteration algorithm. To underline their similarity to regular policies, we call them pseudo-policies.
4 Optimal policy in action space
Consider an action $b$ and a policy $\pi$ such that $b\notin\pi$, assuming without loss of generality that $b$ is an action on the state 1. What is the dot product of the action vector $b_{+}$ and the policy vector $V^{\pi}_{+}$? Mathematically, it is:
$b_{+}V^{\pi}_{+}=r^{b}+(\gamma p^{b}_{1}-1)V^{\pi}_{1}+\dots+\gamma p^{b}_{n}V^{\pi}_{n}=r^{b}+\gamma\sum_{i=1}^{n}p^{b}_{i}V^{\pi}_{i}-V^{\pi}_{1}.$
This quantity is often used in the reinforcement learning literature and is known as an advantage or an appeal of action $b$ with respect to policy $\pi$, denoted as $\textrm{adv}(b,\pi)$ (in this paper we stick to the term advantage). It has an intuitive meaning: the quantity shows the potential gain (or loss) in the infinite sum of discounted rewards if, instead of following the policy $\pi$ from State 1, the agent does the action $b$ and then follows $\pi$.
Advantages are key quantities in MDPs since they define the dynamics of the main MDP algorithms:
  • Given a policy $\pi_{t}$, Policy Iteration algorithm chooses the action $a_{s}$ in a state $s$ with the highest advantage among other actions in $s$.
  • Similarly, while choosing actions to update values, Value Iteration algorithm chooses actions with the highest advantage with respect to current values.
Now let’s establish the relation between an action advantage and a policy hyperplane in the action space. Since the advantage of an action $b$ with respect to a policy $\pi$ is the inner product of the vector $b_{+}$ with the vector $V^{\pi}_{+}$, which is orthogonal to the hyperplane $\mathcal{H}_{\pi}$, a positive advantage means that the vector $b_{+}$ lies above $\mathcal{H}_{\pi}$, and a negative advantage means that $b_{+}$ lies below $\mathcal{H}_{\pi}$. With this observation, we are ready to present the second important result of this paper.
Theorem 4.1.
The problem of MDP solving, or finding a stationary optimal policy $\pi^{*}$ in an MDP, is equivalent to the problem of constructing a hyperplane $\mathcal{H}$ such that at least one action from every state is incidental to it and no action lies above it.
Proof.
$\implies$: Suppose we know the optimal policy $\pi^{*}$. Every action has 00 or negative advantage with respect to it; therefore, none of the action vectors lie above the hyperplane $\mathcal{H}_{\pi^{*}}$. On the other hand, each state has the optimal action, and the corresponding action vector is incidental to $\mathcal{H}_{\pi^{*}}$.
[…]
$a_{i}=\operatorname*{arg\,max}_{b}\textrm{adv}(b,\pi).$
Thus, policy $\pi$ is stationary with respect to greedy upgrade and, consequently, optimal. ∎
The theorem is sufficient to design an algorithm that solves an MDP without ever computing policy values by fitting the highest hyperplane, but we propose a different value-free algorithm later.
5 MDP transformation and affine equivalence
If we try to analyze the dynamics of the policy iteration algorithm in a 2-state MDP (which we do later in Section 6.1 and Figure 3), we see that it is not very convenient to analyze when the rewards for two states are not comparable (e.g., big rewards in one state and small rewards in another state), making all policy hyperplanes steep. Can we modify the MDP to make it flatter? It turns out we can! In this section, we describe the transformation $\mathcal{L}$ of an MDP which changes the values of every policy in one particular state, while keeping all advantages unchanged.
Choose a state $s$ and a real number $\delta\in\mathbb{R}$ — the desired increment of the state value $V^{\pi}(s)$ of every policy $\pi$. We denote this transformation as $\mathcal{L}_{s}^{\delta}$. To perform it, we need to modify every reward in the MDP using the following rule:
  • For every action $a,\mathrm{st}(a)=s$ the reward $r^{a}$ should be modified as:
    $r^{a}:=r^{a}-\delta(\gamma p^{a}_{s}-1),$
  • For every action $b,\mathrm{st}(b)\neq s$ the reward $r^{b}$ should be modified as:
    $r^{b}:=r^{b}-\delta\gamma p^{b}_{s},$
where $p^{a}_{s}$ and $p^{b}_{s}$ are the probabilities of getting to state $s$ from the state of the action.
The geometric interpretation of $\mathcal{L}_{s}^{\delta}$ is as follows. We transform the whole action space with the linear transformation
$(x_{0},x_{1},\ldots,x_{n})\mapsto(x_{0}-\delta x_{s},x_{1},\ldots,x_{n}).$
This transformation maps the designated action zones for each state to itself. Any point on a line $L_{s^{\prime}}$, $s^{\prime}\neq s$, is fixed by this transformation, and any point on the line $L_{s}$ is lifted by $\delta(1-\gamma)$. Coupled with Theorem 3.3, we see that the values of any hyperplane $\mathcal{H}$ stay the same for $s^{\prime}\neq s$, and increases by $\delta$ on the state $s$.
Next, we prove that the advantage of any action with respect to any policy does not change after the transformation. The first case is $\mathrm{st}(a)=s$. Denote the original value vector of a policy as $V_{+}$, the modified value vector as $V^{\prime}_{+}$, the original action vector as $a_{+}$, and the modified action vector as $a^{\prime}_{+}$. Then
$a_{+}V_{+}=\,$ $r^{a}+(\gamma p^{a}_{s}-1)V^{s}_{+}+\gamma\sum_{s^{\prime}\neq s}p^{s^{\prime}}_{a}V^{s^{\prime}}_{+}=$
$r^{a}-\delta(\gamma p^{a}_{s}-1)+(\gamma p^{a}_{s}-1)(V^{s}_{+}+\delta)+\gamma\sum_{s^{\prime}\neq s}p^{s^{\prime}}_{a}V^{s^{\prime}}_{+}=a^{\prime}_{+}V^{\prime}_{+}.$
For action $\mathrm{st}(b)\neq s$ the proof looks essentially the same. Denoting the corresponding action vectors as $b_{+}$ and $b_{+}^{\prime}$ we have:
$b_{+}V_{+}=\,$ $r^{b}+\gamma p^{b}_{s}V^{s}_{+}+\gamma\sum_{s^{\prime}\neq s}p^{s^{\prime}}_{b}V^{s^{\prime}}_{+}=$
$r^{b}-\delta\gamma p^{b}_{s}+\gamma p^{b}_{s}(V^{s}_{+}+\delta)+\gamma\sum_{s^{\prime}\neq s}p^{s^{\prime}}_{b}V^{s^{\prime}}_{+}=b_{+}^{\prime}V_{+}^{\prime}.$
The transformation $\mathcal{L}_{s}^{\delta}$ has a few important properties. For MDPs $\mathcal{M}$ and MDP $\mathcal{M}^{\prime}$, we write $\mathcal{M}^{\prime}=\mathcal{L}_{s}^{\delta}\mathcal{M}$ if $\mathcal{L}_{s}^{\delta}$ transforms $\mathcal{M}$ to $\mathcal{M}^{\prime}$. Note that $\mathcal{M}$ and $\mathcal{M}^{\prime}$ have the same sets of actions and, therefore, the policies are drawn from the same sets of actions. Then, we can make the following observations regarding the dynamics of the policy iteration and value iteration algorithms in MDPs $\mathcal{M}$ and $\mathcal{M}^{\prime}$:
  • Each advantage $\textrm{adv}(a,\pi)$ remains the same, which leads to the same action choices during the policy improvement steps of the policy iteration algorithm in $\mathcal{M}$ and $\mathcal{M}^{\prime}$. Therefore, the transformation $\mathcal{L}_{s}^{\delta}$ preserves the dynamics of the Policy Iteration algorithm.
  • Assume the starting values of the Value Iteration algorithm are $V_{0}$ in $\mathcal{M}$ and $V_{0}^{\prime}$ in $\mathcal{M}^{\prime}$, such that $V_{0}(s)+\delta=V_{0}^{\prime}(s)$. Then, due to preservation of the advantages, the choice of action remains the same at every step of the algorithm. At the same time, since we increase the value of state $s$ by $\delta$ for every pseudo-policy, the difference between the two value vectors remains unchanged. This, in turn, implies that $\mathcal{L}_{s}^{\delta}$ preserves the error vectors $e_{t}=V_{t}-V^{*}$, which describe the dynamics of the value iteration algorithm and define $\epsilon$-optimal policies. It also preserves the values of $V_{t+1}-V_{t}$, the vector used to compute the stopping criteria in the classical Value Iteration algorithm. All these properties together imply that the transformation $\mathcal{L}_{s}^{\delta}$ preserves the dynamics of the Value Iteration algorithm if the initial values are adjusted. We provide more comments on the Value Iteration dynamics in section 6.2.
The dynamics of the main algorithms imply a similarity between MDPs $\mathcal{M}$ and $\mathcal{M}^{\prime}$, based on which we give the following definition:
Definition 5.1.
MDPs $\mathcal{M}$ and $\mathcal{M}^{\prime}$ are affine equivalent if there exists a sequence of operations $\mathcal{L}_{s_{1}}^{\delta_{1}},\mathcal{L}_{s_{2}}^{\delta_{2}},\dots,\mathcal{L}_{s_{k}}^{\delta_{k}}$ such that
$\mathcal{M}^{\prime}=\mathcal{L}_{s_{1}}^{\delta_{1}}(\mathcal{L}_{s_{2}}^{\delta_{2}}(\dots(\mathcal{L}_{s_{k}}^{\delta_{k}}\mathcal{M})\dots)).$
Affine equivalence is a transitive relation on the set of MDPs. An affine equivalence class of MDPs is a maximal subset of MDPs in which every two MDPs $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are affine equivalent.
It can be shown that the order in which transformations are applied does not affect the result. Since transformations are typically applied to all states simultaneously, it is useful to introduce the following notation:
$\mathcal{M}^{\prime}=\mathcal{L}^{\Delta}\mathcal{M}=\mathcal{L}_{s_{1}}^{\delta_{1}}(\mathcal{L}_{s_{2}}^{\delta_{2}}(\dots(\mathcal{L}_{s_{k}}^{\delta_{k}}\mathcal{M})\dots)),$
where $\Delta=(\delta_{1},\delta_{2},\dots,\delta_{n})^{T}$ is a column vector, and the $i$th entry of it, $\delta_{i}$, is the quantity added to every value in state $s_{i}$.
Then, choose some MDP $\mathcal{M}$ and compute its optimal values $(V^{*}(1),\dots,V^{*}(n))$. Apply the following sequence of transformations to $\mathcal{M}$:
$\mathcal{M}^{*}=\mathcal{L}^{\Delta^{*}}\mathcal{M},$ (3)
where $\Delta^{*}=(-V^{*}(1),\dots,-V^{*}(n))^{T}$. The values of the optimal policy $\pi^{*}$ in $\mathcal{M}^{*}$ are 0 in all states. At the same time, since the transformations $\mathcal{L}$ preserve advantages, the advantages of the modified actions $a^{\prime}\in\pi^{*}$ remain 00, which implies that
adv($\textrm{adv}(a^{\prime},\pi^{*})=r^{a^{\prime}}+\sum_{i=1}^{n}\gamma p^{a}_{i}V^{\pi^{\prime}}=r^{a^{\prime}}=0,$
meaning that every action participating in the optimal policy in $\mathcal{M}^{*}$ has reward 00, while every action not participating in the optimal policy has a reward equal to its advantage with respect to the optimal policy, which is always negative. Geometrically, the optimal policy hyperplane $\mathcal{H}_{\pi^{*}}$ is given by the equation $x_{0}=0$, all the action vectors from the optimal policy lie on the hyperplane, while every action not participating in the optimal policy lie below it. In this normal form, the reward of an action becomes a single measure of the goodness of an action.
This picture is convenient to analyze, so we give this special kind of MDPs a definition:
Definition 5.2.
We call an MDP $\mathcal{M}^{*}$ normal if all values of the optimal policy of $\mathcal{M}^{*}$ are equal to zero.
The normalization $\mathcal{M}^{*}$ of an MDP $\mathcal{M}$ is the result of applying the sequence of transformations $\mathcal{L}_{1}^{-V^{*}(1)},\mathcal{L}_{2}^{-V^{*}(2)},\dots,\mathcal{L}_{n}^{-V^{*}(n)}$ to the MDP $\mathcal{M}$ with the optimal policy values $V^{*}$. The normalization is always a normal MDP.
There is exactly one normal MDP in every affine equivalence class. In the next section, we use the normal MDP to analyze the dynamics of the main RL algorithms.
6 Dynamics of main RL algorithms
In this section, we demonstrate the geometrical dynamics of the main MDP algorithms —Value Iteration and Policy Iteration —in the action space. We analyze the dynamics of the normalized MDP since every MDP can be transformed into its normal form while keeping the same affine equivalence class.
6.1 Policy Iteration
Figure 3: Illustration of Policy Iteration dynamics in 2-state MDP.
The example dynamics of the policy iteration algorithm is presented in Figure 3. Policy iteration consists of two steps applied sequentially: policy evaluation and policy improvement. As follows from Proposition 3.2, the policy evaluation step is identical to hyperplane construction. For the policy improvement step, at each iteration, it simply chooses the highest actions with respect to the current policy $\pi_{t}$. Therefore, with every iteration policy iteration algorithm gets a higher hyperplane until the optimal one is reached.
6.2 Value Iteration
Figure 4: Illustration of value iteration dynamics in 2 states MDP.
The dynamics of the Value Iteration algorithm is a little bit more complex and is presented in Figure 4. At each step, value iteration performs the following update:
$V_{t+1}(s)$ $=\max_{a}\left[r(s,a)+\gamma\sum_{s^{\prime}}P(s^{\prime}|s,a)V_{t}(s^{\prime})\right]$
$=V_{t}(s)+r(s,a^{*})+(\gamma P(s|s,a^{*})-1)V_{t}(s)+\gamma\sum_{s^{\prime}\neq s}P(s^{\prime}|s,a^{*})V_{t}(s^{\prime})$
$=V_{t}(s)+\textrm{adv}(a^{*},V_{t}).$
Having this interpretation, it is easy to graphically demonstrate an additional convergence factor described in Mustafin et al. (2024): assume that all current errors $e_{t}=V_{t}-V^{*}$ are positive (in a normalized MDP $e_{t}=V_{t}$). Let’s take a state $s$ with the highest error $e_{t}(s)$. The advantage of the action $a^{*}$ in state $s$, which maximizes advantage with respect to pseudo-policy $\pi_{t}$ (which has values $V_{t}$), is negative. At the same time, its absolute value would be exactly $(1-\gamma)V_{t}(s)$ only if $a^{*}$ is an action participating in the optimal policy and lying exactly on the edge, i.e. this action is a self-loop action. If $a^{*}$ does not participate in the optimal policy, its advantage is at least $(1-\gamma)a^{*}_{+}V_{t+}+\delta^{\prime}$, where $\delta^{\prime}$ is the minimum disadvantage of a non-optimal action. If $a^{*}$ does participate in the optimal policy, but it is not a self-loop action, it lies “inside” its state zone and the absolute value of its advantage is again higher than $(1-\gamma)a^{*}+{V_{t}}_{+}$. These two factors guarantee geometric convergence with a rate faster than $\gamma$.
Geometrically, each iteration of the Value Iteration algorithm not only brings the pseudo-policy implied by the current values closer to optimal policy, but also skews its slope closer to a horizontal. This extra transformation enables achieving $\epsilon$-optimal policy when errors $e_{t}=V_{t}-V^{*}$ are still significantly high. Please refer to Figure 4 and the original paper (Mustafin et al., 2024) for details.
7 Reward balancing solvers
There is one more nice property of a normalized form of a MDP that we did not mention in Section 5: it is trivially solvable! To solve it, one just needs to pick the action with a reward of 0 at every state, and the resulting policy is optimal. This policy has a value of 0 at every state, while any other action has a negative reward, thus their inclusion to a policy will result in negative values.
This is an intriguing property of an MDP. Imagine one needs to solve an MDP $\mathcal{M}$. Suppose that $\mathcal{M}$ can be solved by Policy Iteration starting from a policy $\pi_{0}$ in $T$ steps. If the same algorithm is run on the normalized MDP $\mathcal{M}^{*}$, it will take exactly the same number of iterations $\pi_{0}$ to solve it as it is shown in Section 6.1. However, there is no need to run the algorithm on the normalized MDP: we can just pick the optimal actions with rewards 00 to form the optimal policy. This property of the normalized MDP inspires a new class of solving algorithms that do not compute state values but instead solve the MDP by directly manipulating its rewards. Due to this characteristic, we refer to these value-free algorithms as reward balancing solvers.
Given a random MDP $\mathcal{M}$, it is not possible to normalize it without solving it, because the optimal values are not known. However, we still have the transformations $\mathcal{L}_{s}^{\delta}$ available to us, and we have a goal: if at some point we achieve an MDP in which the maximum action reward in every state is 0, we know that the corresponding actions form the optimal policy. Furthermore, the following lemma sets a stopping criteria for an approximate solution.
Lemma 7.1.
Take the policy $\pi$ in an MDP $\mathcal{M}$ that chooses the actions with the maximum rewards in every state. If the overall maximum reward is 0, and the minimum reward of the actions in the policy $\pi$ is $r_{min}<0$, the policy is $\epsilon$-optimal with $\epsilon=-r_{min}/(1-\gamma)$.
Proof.
We know that values of any policy in this MDP are upper bounded by 0. At the same time, minimum possible values of the policy satisfy $||V^{*}-V^{\pi}||_{\infty}<(0-r_{\min})/(1-\gamma)$. ∎
Having established the stopping criteria, we are now ready to provide an intuition for the solving algorithm. For simplicity, we start by decreasing the reward of every action in the MDP by the overall maximum reward so that the maximum action reward becomes 00 and every action has a non-positive reward. Our goal is then to carefully "flatten" action rewards by applying transformations $\mathcal{L}{s}^{\delta}$ to the states where the highest rewards are below 0, so that the maximum reward in every state becomes closer to 00. The issue is that while increasing the reward of an action $a$ in a state $s$, the transformation $\mathcal{L}{s,\delta}$ also has a negative side effect: it decreases the rewards of all actions in every other state that have a non-zero transition probability to $s$. A straightforward approach would be to account for these side effects, which results in Algorithm 1. The problem with this algorithm lies in its Update step: in fact, the vector $\Delta_{t}$ consists of policy values, which makes this algorithm analogous to the standard Policy Iteration algorithm.
Algorithm 1 Exact Value Free Solver
  Initialize $\mathcal{M}_{0}:=\mathcal{M}$ and $t=1$.
  Policy choice:
  • Select policy $\pi_{t}$ by choosing $a_{s}=\operatorname*{arg\,max}_{a}r^{a},\mathrm{st}(a)=s$ for each state $s$. Denote the reward of the action as $r^{a_{s}}$.
  • if  $r^{a_{s}}=0\quad\forall s$ then
  • Output policy $\pi_{t}$
  • else
  • Update the MDP $\mathcal{M}_{t}=\mathcal{L}^{\Delta_{t}}\mathcal{M}_{t-1}$ by choosing $\Delta_{t}$ such that the values of the policy $\pi_{t}$ in the updated MDP $\mathcal{M}_{t}$ are all 00.
  • Increment $t$ by 1 and return to the policy choice step.
  • end if
However, if we do not aim to produce an exact solution and instead aim for an $\epsilon$-optimal policy, we can develop a novel algorithm, which is not similar to the Value-based approximate algorithm — Value Iteration. The easiest way to simplify the algorithm is to disregard the aforementioned "side" effects, which leads to Algorithm 2. This algorithm performs “safe" updates every iteration: action rewards, which start as negative after the initial normalization, remain negative throughout the run of the algorithm. This is guaranteed by the choice of deltas $\delta_{s}=\max_{a}r_{a}/(1-\gamma p^{a}_{s})$, which is justified from both classic MDP and geometric perspectives. From the classic MDP point of view, the value of every policy at state $s$ is guaranteed to be at least the sum of discounted rewards collected before leaving state $s$. Choosing the maximum among these values over all actions in state $s$ guarantees that the optimal policy has negative values in the updated MDP, which is shown in the following proposition:
Proposition 7.2.
$r_{t}^{a}\leq 0$ for all times $t$ and actions $a$.
Proof.
Recall that we assumed this to hold at step zero. Now we will argue that this property is preserved by applying each map $L_{i}^{\delta_{i}}$ computed in the execution of the VFS. Indeed, in applying this map have
$\delta_{i}=-\max_{a|{\rm st}(a)=i}\frac{r^{a}}{1-\gamma p_{a}^{i}}.$
Since the numerator is negative and denominator is positive, we have
$\delta_{i}\geq 0.$
This immediately implies that for actions $b$ with ${\rm st}(b)\neq i$, the update performed by $\mathcal{L}_{i}^{\delta_{i}}$,
$r^{b}\leftarrow r^{b}-\delta_{i}\gamma p_{b}^{i}$
leaves $r_{b}$ non-positive.
As for actions $a$ with ${\rm st}(a)=i$, we have that the update has the property that
$r^{a}$$\leftarrow$ $r^{a}-\delta_{i}(\gamma p_{s}^{a}-1)$
$=$ $r^{a}+\delta_{i}(1-\gamma p_{s}^{a})$
$\leq$ $r^{a}+\frac{r^{a}}{\gamma p_{s}^{a}-1}(1-\gamma p_{s}^{a})$
$=$ $0$
From the geometric perspective, performing every update $\mathcal{L}_{s}^{\delta_{s}}$ is, in fact, flattening a hyperplane $\mathcal{H}_{\delta_{s}}$, which corresponds to a pseudo-policy that has value $\delta_{s}$ in state $s$ and values of 00 in other states. By design, every action lies below this policy, which means that after the update, the values of the optimal policy remain negative, while the optimal value in state $s$ is closer to 00.
Algorithm 2 Approximate Value Free solver (VFS)
  Initialize $\mathcal{M}_{0}:=\mathcal{M}$ and $t=1$.
  Update:
  • $\forall s$ compute $\delta_{s}=\max_{a}r_{a}/(1-\gamma p^{a}_{s})$.
  • Update the MDP $\mathcal{M}_{t}=\mathcal{L}^{\Delta_{t}}\mathcal{M}_{t-1}$, where $\Delta_{t}=(-\delta_{1},\dots,-\delta_{n})^{T}$.
  • Store the maximum action reward in each state while performing the update. Set $R^{m}_{t}$ to the minimum of these maxima.
  if  $|R^{m}_{t}|/(1-\gamma)\geq\epsilon$ then
     Increment $t$ by 1 and return to the update step.
  else
     Output policy $\pi$, such that $\pi(s)=\operatorname*{arg\,max}_{a}r(a)$.
  end if
Algorithm 2 has the same iteration complexity as the Value Iteration algorithm, but has different convergence properties. First, Value Iteration convergence depends on the initial value guess, while VFS performance does not depend on the quality of initialization. Second, as we know from Mustafin et al. (2024), the Value Iteration algorithm benefits from information exchange between the nodes. In particular, it does not perform well on tree-structured MDPs, potentially achieving the worst possible convergence rate of $\gamma$. In contrast, tree-structured MDPs are the best case for the VFS algorithm. We can show this for a broader case of hierarchical MDPs.
Definition 7.3.
MDP $\mathcal{M}$ is hierarchical if every state can be attributed to a class $c\in\mathbb{N}$ such that every action leads either back to the state $s$ or to states belong to lower classes.
Theorem 7.4.
Applied to hierarchical MDP $\mathcal{M}$ VFS algorithm is guaranteed to converge to the exact solution in at most $C$ iterations where $C$ is a number of hierarchy classes.
Proof.
The theorem can be proven by induction on the class number. We are going to show that after $t$ iterations of the algorithm $t$ lowest hierarchy classes have maximum reward of 00.
Basis: Lowest class of 1 - leaf nodes. Each of states correspondent to these nodes has only self-loop actions, and after first iteration of the algorithm maximum reward on each node becomes 00. That does not change during the run of the algorithm, since probability to transfer to other nodes is 00 and, consequently, $\delta$’s associated with these states remain 00 during the run of the algorithm, so they won’t affect other states.
Step: By induction hypothesis after $t$ steps of the algorithm $\forall s\in c<=t$, maximum reward on state $s$ is equal to 00, which implies that $\delta_{s}$ computed during the $t+1$ steps is also equal to 00. Now recall that for any action $a,\mathrm{st}(a)=s$ it’s new reward $r^{a}_{t+1}$ can be expressed as:
$r^{a}_{t+1}=$ $r^{a}_{t}-\delta_{s}(\gamma p^{a}_{s}-1)+\gamma\sum_{s^{\prime}\neq s}p_{a}^{s^{\prime}}\delta_{s^{\prime}}=$
$r^{a}_{t}-\delta_{s}(\gamma p^{a}_{s}-1)+\gamma\sum_{s^{\prime}\in c,c\leq t}p_{a}^{s^{\prime}}\delta_{s^{\prime}}+\gamma\sum_{s^{\prime}\in c,c>t}p_{a}^{s^{\prime}}\delta_{s^{\prime}}$
Note that for any action $a$ the last sum term is always 00. Then, if class $c$ of $s$ is less or equal to $t$, $\delta_{s}$ and all $\delta_{s^{\prime}}$ in the second summation term are 00, which means that reward of action $a$ does not change and, consequently, the maximum reward on state $s$ remains 00. In the case when $c=t+1$, both summation terms are equal to 00, and $r^{a}_{t+1}=r^{a}_{t}-\delta_{s}(\gamma p^{a}_{s}-1)$. However the quantity $\delta_{s}$ is chosen such that it maximizes the quantity $r^{a}_{t}/(\gamma p^{a}_{s}-1)$, which implies that $max_{a}r^{a}_{t+1}=1$ for every state $s$ in class $t+1$. This concludes the proof of the induction step and the theorem.
The VFS algorithm has interesting convergence properties not only in hierarchical MDPs, but on general MDPs too, which is shown in the following theorem.
Theorem 7.5.
The policy $\pi_{t}$ is $\epsilon_{t}$-optimal where
$\epsilon_{t}=\alpha^{t}\frac{l}{(1-\beta)(1-\gamma)}r_{\max}.$
Here
$\alpha$ $=$ $\max_{s,\pi}\frac{\gamma-\gamma[P_{\pi}]_{ss}}{1-\gamma[P_{\pi}]_{ss}}\leq\gamma,\quad\beta=\max_{s,\pi}\gamma[P_{\pi}]_{ss}\leq\gamma,\quad l=1-\gamma\min_{s,\pi}[P_{\pi}]_{ss}\leq 1,$
and $r_{\max}=-\min_{s}\max_{a}(r^{a}_{0}|\rm{st}(a)=s)$ negation of the minimum among maximum state rewards.
To prove this theorem, we first need the following lemma.
Lemma 7.6.
Suppose $P$ is a stochastic matrix and $D={\rm diag}(I-\gamma P)$. Then the matrix $M$ defined as
$M=I-D^{-1}(I-\gamma P)$
satisfies
$||M||_{\infty}=\max_{i}\frac{\gamma-\gamma[P]_{ii}}{1-\gamma[P]_{ii}}$
Proof.
Observe that the matrix $M$ has zero diagonal and
$M_{ij}$ $=\frac{\gamma[P]_{ij}}{1-\gamma[P]_{ii}}$ when $\mbox{when }i\neq j.$
It is thus a nonnegative matrix, and we have
$||M||_{\infty}$ $=$ $\max_{i}\frac{\gamma\sum_{j}[P]_{ij}}{1-\gamma[P]_{ii}}=\max_{i}\frac{\gamma(1-[P]_{ii})}{1-\gamma[P]_{ii}}$
We now give the proof of Theorem 7.5
Proof.
From the definition of $\Delta_{t}$, we have that
$\Delta_{t}=-D_{\pi_{t}}^{-1}r_{t}^{\pi_{t}},$ (4)
where $D_{\pi_{t}}$ again is a diagonal matrix of $I-\gamma P^{\pi_{t}}$
By the definition of the update $\mathcal{L}^{\Delta_{t}}$, for the policy $\pi_{t}$ (as well as for any other policy),
$V_{t+1}^{\pi_{t}}=V_{t}^{\pi_{t}}+\Delta_{t}.$
Now plugging in Eq. (4),
$V_{t+1}^{\pi_{t}}=V_{t}^{\pi_{t}}-D_{\pi_{t}}^{-1}r_{t}^{\pi_{t}}.$
Next, using the fact $V^{\pi}=(I-\gamma P^{\pi})^{-1}r^{\pi}$ we obtain
$r_{t+1}^{\pi_{t}}=r_{t}^{\pi_{t}}-(I-\gamma P_{\pi_{t}})D_{\pi_{t}}^{-1}r_{t}^{\pi_{t}}.$
Now multiplying both sides by $-D_{\pi_{t}}^{-1}$:
$-D_{\pi_{t}}^{-1}r_{t+1}^{\pi_{t}}=-D_{\pi_{t}}^{-1}r_{t}^{\pi_{t}}-D_{\pi_{t}}^{-1}(I-\gamma P_{\pi_{t}})(-D_{\pi_{t}}^{-1}r_{t}^{\pi_{t}})$ (5)
At time $t+1$, we again choose the next policy $\pi_{t+1}$ implicitly (via choosing the minimizing actions). We thus have the key inequality:
$\Delta_{t+1}$ $=$ $-D_{\pi_{t+1}}^{-1}r_{t+1}^{\pi_{t+1}}$ (6)
$\leq$ $-D_{\pi_{t}}^{-1}r_{t+1}^{\pi_{t}},$ (7)
where the last equality is true because it is the actions in $\pi_{t+1}$ that achieve the maxima at time $t+1$.
We can now put together Eq. (5) and Eq. (6) and Eq. (7) to obtain
$-D_{\pi_{t+1}}^{-1}r_{t+1}^{\pi_{t+1}}\leq-D_{\pi_{t}}^{-1}r_{t}^{\pi_{t}}-D_{\pi_{t}}^{-1}(I-\gamma P_{\pi_{t}})(-D_{\pi_{t}}^{-1}r_{t}^{\pi_{t}}).$
We rewrite this as
$\Delta_{t+1}$$\leq$ $\Delta_{t}-D_{\pi_{t}}^{-1}(I-\gamma P_{\pi_{t}})\Delta_{t}$
$=$ $M_{\pi_{t}}\Delta_{t},$
where similarly to before
$M_{\pi}:=I-D_{\pi}^{-1}(I-\gamma P_{\pi}).$
Using induction and Lemma 7.6, we obtain
$||\Delta_{t}||_{\infty}\leq\alpha^{t}||\Delta_{0}||_{\infty},$
where $\alpha$ comes from the definition of the theorem.
We have thus shown that the size of the update to the value function decreases exponentially in the infinity norm. We next translate this into a direct suboptimality bound on $V^{\pi_{t}}$.
Starting from
$r_{t}^{\pi_{t}}=-D_{\pi_{t}}\Delta_{t},$
and
$V_{t}^{\pi_{t}}=(I-\gamma P_{\pi_{t}})^{-1}r_{t},$
we put the last two equations together to obtain
$V_{t}^{\pi_{t}}=-(I-\gamma P_{\pi_{t}})^{-1}D_{\pi_{t}}\Delta_{t}$
Let $c=||(I-\gamma P_{\pi_{t}})^{-1}D_{\pi_{t}}||_{\infty}$. Clearly,
$c\leq||(I-\gamma P_{\pi_{t}})^{-1}||_{\infty}||D_{\pi_{t}}||_{\infty}\leq\frac{1}{1-\gamma}l.$
We then have that
$||V_{t}^{\pi_{t}}||_{\infty}\leq c\alpha^{t}||\Delta_{0}||_{\infty}.$
Let us use bound
$||\Delta_{0}||\leq\frac{r_{\max}}{1-\beta},$
so that
$||V_{t}^{\pi_{t}}||_{\infty}\leq c\frac{\alpha^{t}}{1-\beta}r_{\max}.$
The RHS of the above equation is $\epsilon_{t}$ from the definition of the theorem. Since $V_{t}^{\pi}$ is always nonpositive (since rewards are nonpositive throughout execution of the VFS by Proposition 7.2), this infinity norm bound implies that for any policy $\pi$,
$V_{t}^{\pi}\leq V_{t}^{\pi_{t}}+\epsilon_{t}{\bf 1}.$
Since for any policy $\pi$ we also have that
$V_{t}^{\pi}=V_{0}^{\pi}+\Delta_{0}+\cdots+\Delta_{t-1},$
we obtain that
$V_{0}^{\pi}\leq V_{0}^{\pi_{t}}+\epsilon_{t}{\bf 1.}$
We have thus established the $\epsilon_{t}$ sub-optimality of $\pi_{t}$. ∎
The theorem has an interesting corollary which allows the algorithm to converge to exact solution by applying action filtering.
Corollary 7.7.
For all actions $a$,
$\left|r_{t}^{a}-\textrm{adv}(a,\pi^{*})\right|\leq 2r_{\max}\frac{\gamma^{t}}{1-\gamma}.$
Proof.
Let $\pi_{t}$ be the policy defined earlier achieving the minimum at time $t$ and let $\pi^{*}$ be an optimal policy. We have that
$0\geq V_{t}^{\pi^{*}}\geq V_{t}^{\pi_{t}}\geq-r_{\max}\frac{\gamma^{t}}{1-\gamma}.$
Since
$V_{t}^{\pi^{*}}=V_{0}^{\pi^{*}}+\sum_{k=0}^{t-1}\Delta_{k},$
we obtain that if we defube
$d_{t}=V_{0}^{\pi}-\left(-\sum_{k=0}^{t-1}\Delta_{k}\right),$
then
$||d_{t}||_{\infty}\leq r_{\max}\frac{\gamma^{t}}{1-\gamma}.$ (8)
Next, we have
$r_{t}=r_{0}+\sum_{k=0}^{t-1}(I-\gamma P)\Delta_{k}.$
We can write this as
$r_{t}=r_{0}+\sum_{k=0}^{t-1}(\gamma P-I)(V_{0}^{\pi^{*}}-d_{t}),$
or
$r_{t}=r_{0}+\gamma PV_{0}^{\pi^{*}}-V_{0}^{\pi^{*}}-(\gamma P-I)d_{t}.$
Finally, for every action $a$,
$r_{t}^{a}=\textrm{adv}(a,\pi^{*})+(\gamma P-I)d_{t}.$
Finally, using the bound $||\gamma P-I||_{\infty}\leq 2$ and Eq. (8) concludes the proof. ∎
The two theorems in this section suggest that while having the same iteration computational complexity, VFS and VI are two different algorithms. Section 9 demonstrates the experimental comparison of the two algorithms.
8 Algorithm for unknown MDPs.
In the previous section of the paper, we considered a theoretical case of fully known MDPs, where all transition probabilities $\mathcal{P}$ are known and available to the solver. In practice, the assumption of a fully known MDP is usually unfeasible, and the solver only has access to transition data in the form $(s,s^{\prime},a,r^{a})$, which means that after choosing an action $a$ in state $s$, the agent receives a reward $r^{a}$ and transitions to state $s^{\prime}$, sampled from $P(s^{\prime}|a)$. This constraint requires a few adjustments to the algorithm, but before we discuss them, we need to switch from state-based to action-based notation. From now on, the reward vector $r_{t}$ is of length equal to the number of actions $m$ and consists of all available action rewards from every state. The transition matrix $P^{\pi}$ has a size of $m\times m$, and its element $P[i,j]$ is the probability that after choosing action $i$, the next action chosen is action $j$. This value is non-zero only if the agent has a non-zero probability of appearing in state $s^{\prime}=\rm{st}(j)$, and policy $\pi$ chooses action $j$ in $s^{\prime}$. For our analysis, we need an additional vector $R_{t}$, a vector of length $m$, where for every action $a$, it stores the maximum reward available in the state of action $a$ at time $t$, $R_{t}[i]=\max_{b}r^{b},\,\rm{st}(b)=\rm{st}(a)$. For convenience we abuse notation and define $R_{t}(s)=\max_{b}r^{b},\,\rm{st}(b)=s$.
The first adjustment we need to make is to give up on self-loop coefficients and the additional convergence they provide, since self-loop probabilities are no longer known. Therefore, without the self-loop coefficients, but with known transition probabilities, the reward balancing update would look like
$r_{t+1}=r_{t}-R_{t}+\gamma PR_{t};$ (9)
However, we no longer have access to the transition probabilities $P^{\pi_{t}}$, but only to samples from it. Here, we employ a free sampling model, which is common in the literature analyzing the convergence of the Q-learning algorithm. This model assumes that the solver can freely sample any number of transitions for any action at any time. For the stability of the algorithm, during time $t$, we sample not one, but $k$ transitions for every action. These transitions are represented by matrices $P^{i}_{t}$, each of which has a single $1$ and the rest 00 in every row, with the expectation $\operatorname{\mathbb{E}}[P^{i}_{t}]=P$. Then, the algorithm update becomes:
$r_{t+1}=r_{t}-R_{t}+\gamma\left(\frac{P^{1}_{t}+\dots+P^{k}_{t}}{k}\right)R_{t}.$ (10)
We denote the transition estimation matrix above as $P_{t}$:
$P_{t}=\frac{P^{1}_{t}+\dots+P^{k}_{t}}{k},\quad\operatorname{\mathbb{E}}[P_{t}]=P.$
The following theorem analyzes algorithm convergence:
Theorem 8.1.
Choose
$k\lesssim\frac{1}{\epsilon^{2}}\frac{r_{\max}^{2}}{(1-\gamma)^{3}}\log n\log\frac{1}{\epsilon(1-\gamma)},$
where $r_{\max}$ is the same as in Theorem 7.5, and run stochastic reward balancing for
$t\lesssim\frac{1}{1-\gamma}\log\frac{1}{\epsilon(1-\gamma)},$
steps. Let $\pi_{t}$ be the policy where every node takes the action with the maximal reward at time $t$. Then $\pi_{t}$ is $\epsilon$-optimal with constant probability bounded away from zero.
To parse this theorem, note that, ignoring various logs, we do $k\sim(1-\gamma)^{-3}$ samples per step with $t\sim(1-\gamma)^{-1}$ steps total, for a sample complexity of $\sim(1-\gamma)^{-4}$. This matches the state-of-the-art sample complexity of Q-learning (Li et al., 2024). However, unlike Q-learning, this algorithm is perfectly parallelizable, which makes an improvement over the state of the art presented in Woo et al. (2023).
The key quantity to track during the run of the algorithm is a minimum of state maximum rewards, $R^{min}_{t}=\min R_{t}$, since it provides a stopping criteria as show in Lemma 7.1. The following proposition demonstrates convergence of this quantity that during the run of the algorithm.
Proposition 8.2.
With probability one, we have that for all $t$,
$R^{\min}_{t+1}\geq\gamma R^{\min}_{t}$
Proof.
Consider some state $s$ which does not necessary achieve the minimum in the definition of $R^{min}_{t}$. Instead, let us write
$R_{t}(s)=L_{t}+o_{t}(s),$
where $R_{t}(s)$ is the maximum of action rewards in state $S$ and
$o_{t}(s)\geq 0.$
Consider any action action $u$ with ${\rm st}(u)=s$. Then from we have,
$r_{t+1}^{u}$ $=$ $r_{t}^{u}-R_{t}(s)+(\gamma PR_{t})[u]$
$\geq$ $r_{t}^{u}-(R^{\min}_{t}+o_{t}(s))+\gamma R^{\min}_{t},$
where the second equation used the bound
$(PR_{t})[u]\geq R^{\min}_{t},$
which follows by the stochasticity of $P$ and definition of $R^{\min}_{t}$. Taking the max of both sides over actions $u$ in state $s$, we obtain
$R_{t+1}(s)$ $=$ $\max_{u|{\rm st}(u)=s}r_{t+1}^{u}$
$\geq$ $(L_{t}+o_{s}(t))-(L_{t}+o_{s}(t))+\gamma L_{t}(s)$
$=$ $\gamma L_{t}(s)$
Since this is true for every $s$, the proposition is proved. ∎
During the run of the algorithm rewards of all actions are always negative, which is demonstrated in the following proposition, similar to Proposition 7.2
Proposition 8.3.
With probability one, $r_{t}\leq 0$ for all $t$.
Proof.
We will prove this my induction. This holds at time zero by the inductive assumption. Now suppose it holds at time $t$. In particular, $\max r_{t}\leq 0$. It suffices to prove that $\max r_{t+1}\leq 0$. But:
$\max r_{t+1}$ $=$ $\max(r_{t}-R_{t}+\gamma P_{t}R_{t})$
$\leq$ $R_{r}-R_{r}+\max\gamma P_{t}R_{t}$
$=$ $\gamma\max P_{t}R_{t}.$
By the inductive hypothesis we have $\max r_{t}\leq 0$ and combining that with the stochasticity of $P_{t}$, we obtain the RHS is nonpositive. This completes the proof. ∎
Corollary 8.4.
With probability one,
$||\max r_{t}||_{\infty}\leq r_{\max}.$
Proof.
By Proposition 8.3, we have that $r_{t}$ is a nonpositive vector, so that
$||\max r_{t}||_{\infty}=|\min_{s}~{}~{}R_{t}(s)|$
By further applying Proposition 8.2, we obtain that the RHS is nonincreasing. Since initially the RHS is upper bonded by $r_{\max}$, this completes the proof. ∎
Next, let us rewrite our update rule to have an explicit expression for the stochastic error at each time step. Instead of Eq. 10 let us write
$r_{t+1}=r_{t}-\max r_{t}+\gamma PR_{t}+z_{t},$
where
$\operatorname{\mathbb{E}}[z_{t}|r_{t},\ldots,r_{0}]=0.$
Specifically, we have
$z_{t}=\gamma(P_{t}-P)R_{t}.$
The size of the noises $z_{t}$ plays an important role in our analysis. What turns out to matter is the infinity norm of the sum of these random variables, which is bounded in the next proposition.
Proposition 8.5.
The random variable
$Z_{t}=\sum_{l=0}^{t-1}z_{l}$
is a sub-Gaussian random variable with parameter $\sigma=r_{\max}\sqrt{t/k}$.
Proof.
Observe that each $z_{l}(s)$ is the average of $k$ random variables whose support lies in $[-r_{\max},r_{\max}]$ by Corollary 8.4. Let us denote these random variables as $e_{l}^{i}(s)$, so that
$z_{l}(s)=\frac{e_{l}^{i}(s)}{k}.$
We thus have that
$P(|\sum_{l=0}^{t-1}z_{l}(s)|\geq\epsilon)=P(|\sum_{l=0}^{t-1}\sum_{i=1}^{k}\frac{e_{l}^{i}}{k}|\geq\epsilon)=P(|\sum_{l=0}^{t-1}\sum_{i=1}^{k}e_{l}^{i}|\geq k\epsilon).$
We can apply Azuma’s inequality to the last expression because if we list $e_{l}^{i}$ in order of increasing $l$, then we have a martingale difference sequence. This gives
$P(|\sum_{l=0}^{t-1}z_{l}(s)|\geq\epsilon)\leq 2e^{-k^{2}\epsilon^{2}/(r_{\max}^{2}kt)}=2e^{-k\epsilon^{2}/(r_{\max}^{2}t)}.$
We conclude that $\sum_{l=0}^{t-1}z_{l}(s)$ is a sub-Gaussian random variable with parameter $\sigma=r_{\max}\sqrt{t/k}$.
Proposition 8.6.
$E||Z_{t}||_{\infty}\lesssim r_{\max}\sqrt{\frac{t\log n}{k}}$
Proof.
Using the previous proposition, this now follows from a standard result on the maximum of sub-Gaussian random variables (see exercise 2.5.10 of Vershynin (2018)). ∎
Given a policy $\pi$, let us denote by $G_{t}^{\pi}$ the vector
$G_{t}^{\pi}=Q_{0}^{\pi}-Q_{t}^{\pi},$
where recall $Q_{t}^{\pi}$ is the value vector of the policy $\pi$ under the rewards $r_{t}$. This vector has as many entries as the number of actions. Note that because $r_{t}$ is now a random variable, so is $Q_{t}^{\pi}$ and thus also $G_{t}^{\pi}$.
Lemma 8.7.
$\operatorname{\mathbb{E}}\max_{\pi,\pi^{\prime}}||G_{t}^{\pi}-G_{t}^{\pi^{\prime}}||_{\infty}\lesssim\frac{1}{1-\gamma}E||\sum_{l=0}^{t-1}z_{l}||_{\infty}.$
Proof.
Indeed, for any fixed policy $\pi$,
$Q_{t}^{\pi}=(I-\gamma P_{\pi})^{-1}r_{t},$
where $P_{\pi}$ is the action-to-action transition matrix. Thus
$Q_{t+1}^{\pi}=Q_{t}^{\pi}+(I-\gamma P_{\pi})^{-1}(\gamma P-I)R_{t}+(I-\gamma P_{\pi})^{-1}z_{t}.$
A key observation is that the second term on the RHS does not depend on the policy, despite appearing to do so. Indeed, policy $\pi_{t}$ specifies which action we would choose after reaching a certain state, but all actions of one state $s$ has the same reward, therefore
$PR_{t}=P_{\pi}R_{t},$
for any policy $\pi$. We thus have
$Q_{t+1}^{\pi}=Q_{t}^{\pi}-R_{t}+(I-\gamma P_{\pi})^{-1}z_{t}.$
The independence of $\pi$ of the second term on the RHS is clear here in the above equation.
Now, iterating this recursion, we obtain that for any two policies $\pi,\pi^{\prime}$,
$G_{t}^{\pi}-G_{t}^{\pi^{\prime}}=(I-\gamma P_{\pi})^{-1}\sum_{l=0}^{t-1}z_{l}-(I-\gamma P_{\pi^{\prime}})^{-1}\sum_{l=0}^{t-1}z_{l}.$ (11)
Thus
$||G_{t}^{\pi}-G_{t}^{\pi^{\prime}}||_{\infty}\leq||(I-\gamma P_{\pi})^{-1}-(I-\gamma P_{\pi^{\prime}})^{-1}||_{\infty}||\sum_{l=0}^{t-1}z_{l}||_{\infty}.$
Therefore
$||G_{t}^{\pi}-G_{t}^{\pi^{\prime}}||_{\infty}\leq\frac{2}{1-\gamma}||\sum_{l=0}^{t-1}z_{l}||_{\infty}.$
Since this bound holds for any policy $\pi$ and any policy $\pi^{\prime}$, we have in fact proven
$\max_{\pi,\pi^{\prime}}||G_{t}^{\pi}-G_{t}^{\pi^{\prime}}||_{\infty}\leq\frac{2}{1-\gamma}||\sum_{l=0}^{t-1}z_{l}||_{\infty}.$
We can now prove our main result of this section.
Proof of Theorem 8.1.
By Proposition 8.2, at $t=t_{\rm total}$ we have that
$R^{\min}_{t}\geq-\epsilon(1-\gamma).$
It follows that $Q_{t}^{\pi_{t}}\geq-\epsilon{\bf 1}$. On the other hand, using nonpositivity of rewards from Proposition 8.3, we obtain that $Q_{t}^{\pi}\leq 0$ for any policy $\pi$. We conclude that the policy $\pi_{t}$ is $\epsilon$-optimal under the rewards $r_{t}$.
However, what we need to bound is the optimality of $\pi_{t}$ under the initial rewards $r_{0}$. Observe that for any policy $\pi$ we have by definition of $G_{t}^{\pi}$,
$Q_{t}^{\pi}=Q_{0}^{\pi}+G_{t}^{\pi}.$
By $\epsilon$-optimality of $\pi_{t}$ under the rewards $r_{t}$, we have that for any policy $\pi$,
$Q_{0}^{\pi}+G_{t}^{\pi}\leq Q_{0}^{\pi_{t}}+G_{t}^{\pi_{t}}+\epsilon{\bf 1},$
or
$Q_{0}^{\pi}-Q_{0}^{\pi_{t}}\leq G_{t}^{\pi_{t}}-G_{t}^{\pi}+\epsilon{\bf 1}.$ (12)
We conclude that the event that the policy $\pi_{t}$ is not $2\epsilon$ optimal under $r_{0}$ can occur only if for some policy $\pi$, we have
$||G_{t}^{\pi}-G_{t}^{\pi_{t}}||_{\infty}\geq\epsilon.$
Let us call that event $\mathcal{A}$.
We can thus bound
$P(\mathcal{A})$$\leq$ $P(\max_{\pi,\pi^{\prime}}||G_{t}^{\pi}-G_{t}^{\pi^{\prime}}||_{\infty}\geq\epsilon)$
$\leq$ $\frac{E\max_{\pi,\pi^{\prime}}||G_{t}^{\pi}-G_{t}^{\pi^{\prime}}||_{\infty}}{\epsilon}$
$\leq$ $\frac{1}{1-\gamma}\frac{E||\sum_{l=0}^{t-1}z_{l}||}{\epsilon},$
where the second inequality is just Markov’s inequality while the third inequality uses Lemma 8.7. Now using Proposition 8.6, we obtain
$P(\mathcal{A})\lesssim\frac{r_{\max}}{1-\gamma}\sqrt{\frac{t\log n}{k}}\frac{1}{\epsilon},$
and the choice of $k$ in the theorem statement makes this bounded away from $1$. This concludes the proof (after rescaling $\epsilon\leftarrow\epsilon/2$ so that we are considering whether $\pi_{t}$ is $\epsilon$-optimal under $r_{0}$).
9 Experimental results
In the section we present the performance results of the VFS Algorithm 2 on three experiments. We compare the convergence of the VFS algorithm with the Value Iteration algorithm with different learning rates. The comparison is done on three kinds of MDPs:
  • Random MDP: MDP with 10 states, 1 to 3 actions on every state, and a random reward.
  • Grid World MDP: 100 states correspond to the cells in a 10 by 10 grid. Possible actions are “Up”, “Left”, “Down” and “Right”, which are available at any state unless they bring the agent behind the border of the grid. Rewards for each action are equal to the sum of the cell coordinates plus a small random component.
  • Cycle MDP: numbered states organized into a cycle. Each state has 3 available actions, which move the agent to the first, second and third state forward. Reward is equal to the number of the state plus small random component.
Additionally to the parameters above, all the MDPs are created with an extra “execution probability” parameter. Each action’s transition probabilities are multiplied by the execution probability, and the remainder probability is assigned as the probability to stay in the same state (self-loop).
Our results are presented in Figure 5. Subfigures show that when execution probability is set to $1$, VI and VFS algorithms have similar performance, but then as we decrease it, the performance of VFS improves and the performance of the VI declines.
(a) Random MDP
(b) Grid World
(c) Cycle
(d) Legend
Figure 5: Experimental results for 3 kind of MDPs. X-axis: action execution probability. Y-axis: average number of algorithm iterations before convergence to $\epsilon$. Green colors - VFS algorithm, Red colors- VI algorithms. Circle marker with light color - $0.5$ learning rate, Square marker with medium color - $0.75$ learning rate, Upper triangle marker with dark color - $1.0$ learning rate.
10 Discussion and Concluding remarks
In this paper, we have presented a new geometric interpretation of Markov Decision Processes, establishing the equivalence between MDP problems and specific geometrical problems. Based on this interpretation, we developed new algorithms for both known and unknown MDP cases, showing state-of-the-art convergence properties in some cases and improving on the state-of-the-art in others. At the same time, the authors believe that these are only the first results provided by this new interpretation. This work invites researchers in the field to apply the new perspective on MDPs to their research, whether in the analysis of existing algorithms or the development of new ones.